OpenStudy (anonymous):

Assume that you have a car worth $7300 and you wish to insure it for its full replacement value if it is stolen. If there is a 2% chance that the car will be stolen, what would a fair premium price be? A. $133.94 B. $152.98 C. $148.98 D. $161.30

jimthompson5910 (jim_thompson5910):

P(Stolen)*(Gain) + P(Not stolen)*(Gain) = 0 0.02*(-7300) + (1-0.02)x = 0 -146 + 0.98x = 0 0.98x = 146 x = 146/0.98 x = 148.97959 which to the nearest penny rounds to x = 148.98 So a fair premium price would be $148.98